# AI-boosted rare event sampling to characterize extreme weather

### v1

@@ METHODS
S4 Choosing the hyperparameters of the algorithm
@@
In the Diffusion Monte Carlo (DMC) method, the splitting functions $\left(V_{k}\right)_{k}$ can critically influence the behavior of the algorithm. A frequently used choice is $V_{k}(X)=C_{k}\theta(X)$, where $C_{k}>0$ controls the duplication rate of walkers. While simple, this formulation has drawbacks. In nonlinear systems, walkers with large $\theta$ values may produce an excessive number of offspring, eventually leading to the so-called extinction phenomenon in which the population becomes overly concentrated around a few trajectories. Although such a scheme remains unbiased, the variance of the resulting estimator can be prohibitively large. On the opposite end, if selection is too weak, particle weights are nearly equal, and the method essentially reduces to naive Monte Carlo sampling—again yielding high variance. Proper tuning is therefore essential to balance these two extremes.
To alleviate these issues, [19] proposed a modification of DMC designed to improve robustness. The idea is to rescale the score function $\theta$ dynamically at each resampling step so that its distribution matches a prescribed target distribution $\nu_{k}$, most often taken as Gaussian. The transformed score, denoted $\theta_{k}^{\prime}$, is then used in the splitting and pruning procedure. This approach is known as Quantile Diffusion Monte Carlo (QDMC).
@@
In our experiments, we observed that when the number of walkers NN is too small, the quantile-mapping step—based on the empirical estimate of $F_{\theta_{k}}$ which only relies on the ranking between walkers—can have an undesired effect, sometimes assigning very different weights to walkers with very close scores. To avoid this issue, we opted for a simpler normalization procedure described in Eq. (5) in the main text.
S4.2 Choice of the resampling scheme
@@
S4.3 Scaling the constants $C_{k}$
For the splitting functions, we chose to set $V_{k}=C_{k}\theta_{k}$, where $C_{k}>0$ is a constant to tune. A higher value of $C_{k}$ results in a larger number of walker clones, making the algorithm more selective. However, if $C_{k}$ is too large, it may lead to the extinction phenomenon described in Section S4.1, whereas if it is too small, the algorithm behaves similarly to direct numerical sampling by not sampling enough extreme events. It is often desirable not to use the same constant $C_{k}$ for all resampling steps. In our case, we are forecasting the future state of the trajectory to compute the score function, and the uncertainty in the forecast decreases as we approach the final time $t_{f}$. Therefore, we chose to set $C_{k}=Ce^{-\alpha(t_{f}-t_{k})}$, where $C>0$ is a constant and $\alpha>0$ is a decay factor. A similar approach is used in [19]. Both $C$ and $\alpha$ are hyperparameters of the algorithm. Using a non-exhaustive grid search, we found that setting $C=2.0$ and $\alpha=0.02$ yielded good results in terms of variance of the final estimators of rare event probabilities. Furthermore, we turned off resampling at the first three resampling times, $t_{1}$, $t_{2}$, and $t_{3}$, by setting $C_{1}=C_{2}=C_{3}=0$ to ensure that the walkers were sufficiently separated in phase space before actually applying a resampling.
@@
S5 Computing error bars for the DMC algorithm
To compute variance estimates with the DMC algorithm, we performed 10 independent runs of the algorithm with the same hyperparameters and the same initial conditions. We then computed the empirical variance of the return period estimates across the 10 runs. The same methodology is used in [22, 18, 29].
@@
The intuition behind these two forms is the following: the optimistic estimator considers each trajectory as an independent contribution, similar to importance sampling [69], while the pessimistic estimator aggregates all descendants of a given ancestor into a single effective data point. In practice, the effective number of independent samples lies between these two limiting cases. As there is no theoretical guarantee that one of these estimators is better than the other, we preferred to use the empirical variance computed from the 10 independent runs of the algorithm to compute variance speed-up factors (see Eq. (13) and Fig. 3 in the main text). In Fig. S2, we verify that the empirical variance lies between the pessimistic and optimistic estimators, within statistical error.
@@
FIG. S2: Comparison of speed-up factors obtained via the empirical, the pessimistic (Eq. (S3)) and optimistic (Eq. (S4)) variance estimates. The experiments shown here are described in Extended Data Table II in the main text, for Chicago (left panel) and France (right panel). The variance speed-up factor measures the reduction in the variance of return time estimates compared to direct numerical sampling (Eq. (13) in the main text). The blue curve shows the empirical speed-up ratio obtained by running the AI+RES algorithm 10 times and computing the variance of the return time estimates across the 10 runs. The green curve shows the mean over the 10 optimistic variance estimates, each computed from a single realization of the algorithm. The orange curve shows the mean over the 10 pessimistic variance estimates, each computed from a single realization of the algorithm.
S6 Interpretation of the return times
@@
FIG. S3: Most Recent Common Ancestor Distance (MRCAD) for AI+RES sampled trajectories. The figure shows results for France and Chicago with the setup given in Extended Data Table II in the main text. The MRCAD is the average number of resampling steps since the current population last shared a common ancestor (max = 7, min = 1, with $K=6$ resampling steps). The solid lines show the mean MRCAD over 10 independent realizations of the AI+RES algorithm, with the shading indicating the $95\%$ confidence interval. A larger MRCAD indicates greater trajectory diversity. The AI+RES generated trajectories maintain substantial diversity, yielding a representative catalog of rare events.
@@
S8 Using PLASIM itself to produce forecasts
@@
FIG. S4: Return-time curve comparison between AI+RES and PFS+RES. The setup matches that described in Extended Data Table II in the main text, except that it uses $N=200$ walkers (instead of $N=400$), since larger values are computationally prohibitive for PFS+RES. Black dots are the empirical return periods obtained with a $50,000$-member control simulation. The red dots are from this control simulation but using the same computational budget as the AI+RES algorithm ($N=200$). The solid blue line shows the median return time curve produced by 10 independent realizations of the AI+RES algorithm with $N=200$ walkers and the blue-shaded area represents the 10th to 90th percentile range. The solid green line shows the median return time curve produced by 10 independent realizations of the PFS+RES algorithm with $N=200$ walkers and the green shaded area represents the 10th to 90th percentile range. The gray shaded area is obtained by fitting different GDP distributions with independent $N=200$ training datasets, and showing the 10th to 90th percentile range.
@@
The Perfect-Forecast-System+RES (PFS+RES) approach is designed to provide an upper bound on algorithmic performance. As noted earlier, using the AI-emulator introduces an imperfect approximation of the score function due to the emulator’s limited weather forecast skill, particularly beyond 10–15 days of lead time. Here, we propose performing the ensemble forecast directly with the PlaSim GCM, thereby obtaining a near-perfect forecast to use in the score function.
@@
We conducted experiments using the same setup as the AI+RES experiments, with parameters listed in Extended Data Table II in the main text. The only difference is that the number of walkers was set to $N=200$ instead of $N=400$, as the latter is computationally prohibitive for PFS+RES.
@@
Extended Data Fig. 7 of the main text shows the return-time curves for AI+RES and PFS+RES, demonstrating that PFS+RES can produce highly accurate return-period estimates with smaller variance than AI+RES. To quantify precisely how much we lose in terms of variance when moving from a perfect forecast system (PFS+RES) to an imperfect one (AI+RES), we compared the speed-up factors relative to DNS for both methods in Extended Data Fig. 7. Importantly, these speed-up factors assume the cost of running ensemble forecasts is negligible. While this is not the case for PFS+RES, the comparison provides an indication of the potential performance achievable with near-perfect forecasts, under the assumption that forecast computations are inexpensive relative to the cost of running the physical model (here, PlaSim).
S9 Additional results
@@
FIG. S6: Correlation between predicted and actual $A_{L}(t_{f})$ across forecast systems and regions. The blue curves correspond to forecasts made with the AI emulator within AI+RES for Chicago, while gold shows AI emulator results for France. Red curves represent forecasts made with the PlaSim GCM itself—referred to as the Perfect Forecast System (PFS) for the France region. Results are based on 10 independent realizations of the experiments in Extended Data Table II in the main text. For PFS, the number of walkers is $N=200$ instead of $N=400$ for the AI+RES experiments, and the analysis is limited to France, for computational reasons. At each resampling step in AI+RES, score functions $\theta_{k}^{i}$ are computed from ensemble forecasts (with $M=100$ members) with the AI emulator for each PlaSim walker (Eq. (6) in the main text). Here, three resampling steps are used (at $t_{f}-t_{k}=10$, 55, and 0 days before the event). We show the correlation between predicted (ensemble mean from the emulator) and actual (PlaSim realizations) $A_{L}(t_{f})$ for all events (left) and for events with return periods >100>100 years (right). For rare events, the emulator forecasts are notably less skillful at the earliest resampling time in France compared to Chicago, likely reflecting the stronger role of soil moisture in the former region (since soil moisture is absent from emulator inputs but evolved by PlaSim). This result explains the higher variance of rare-probability estimators in France compared to Chicago (left panel of Fig. 3 in the main text). Similarly, the smaller variances obtained for PFS-RES compared to AI-RES (Extended Data Fig. 7. of the main text), can also be attributed to an overall higher forecast skill of PFS both for typical and extreme events.

### v2

@@ Acknowledgments
S3.4 Ensemble Forecasting
Ensemble forecasts are generated by adding independent initial condition perturbations to all atmospheric variables at the beginning of the forecast. Perturbations are sampled from a Gaussian distribution with mean 0 and a standard deviation such that the resulting ensemble spread over time is similar to that of the PlaSim GCM with our selected magnitude of perturbations to the initial surface pressure (see Section S4.5).
Boundary Variables Land-sea mask, surface roughness, and surface geopotential height (constant); Sea surface temperature, sea ice cover, and total incoming solar radiation (yearly repeating)
Prognostic Atmospheric Variables Specific humidity, temperature, U and V wind, and geopotential height at 50, 100, 150, 200, 250, 300, 400, 500, 600, 700, 850, 925, and 1000 hPa
Prognostic Surface Variables $\log(\text{Surface pressure})$ and air temperature at 2 m
Diagnostic Variables Precipitation accumulated over 6 Hours
Table S2: Description of the variables input to and predicted by the AI emulator. Boundary variables are prescribed variables that are only input to the AI emulator. Prognostic variables refer to the variables both input to and predicted by the AI emulator. Diagnostic variables are only predicted by the AI emulator.
FIG. S1: Weather forecast skill of the AI emulator. Global RMSE (top) and ACC (bottom) as a function of lead time for T2m (left) and Z500 (right). Gold and red lines show the AI emulator with a single-member forecast and a 100-member ensemble forecast, respectively. The blue line corresponds to a persistence forecast. Horizontal black lines in the RMSE panels indicate the climatological forecast. Dashed horizontal lines in the ACC panels at 0.6 mark the conventional threshold below which forecasts are no longer considered skillful.
S4 Rare event algorithm: Diffusion Monte Carlo
We used a rare event algorithm called Diffusion Monte Carlo (DMC), in a version that closely follows the formulations proposed in [14] and [16]. The procedure is described in Algorithm 1.
@@
As mentioned, the splitting functions $\left(V_{k}\right)_{k}$ can critically influence the behavior of the algorithm. A frequently used choice is $V_{k}(X)=C_{k}\theta(X)$, where $C_{k}>0$ controls the duplication rate of walkers. While simple, this formulation has drawbacks. In nonlinear systems, walkers with large $\theta$ values may produce an excessive number of offspring, eventually leading to the so-called extinction phenomenon in which the population becomes overly concentrated around a few trajectories. Although such a scheme remains unbiased, the variance of the resulting estimator can be prohibitively large. On the opposite end, if selection is too weak, particle weights are nearly equal, and the method essentially reduces to naive Monte Carlo sampling—again yielding high variance. Proper tuning is therefore essential to balance these two extremes.
To alleviate these issues, [14] proposed a modification of DMC designed to improve robustness. The idea is to rescale the score function $\theta$ dynamically at each resampling step so that its distribution matches a prescribed target distribution $\nu_{k}$, most often taken as Gaussian. The transformed score, denoted $\theta_{k}^{\prime}$, is then used in the splitting and pruning procedure. This additional step gives the algorithm its full name: Quantile Diffusion Monte Carlo (QDMC).
@@
In our experiments, we observed that when the number of walkers $N$ is too small, the quantile-mapping step—based on the empirical estimate of $F_{\theta_{k}}$, which only relies on the ranking between walkers—can have an undesired effect, sometimes assigning very different weights to walkers with very close scores. To avoid this issue, we opted for a simpler alternative by rescaling the score function $\theta$ at each resampling step using the following transformation
$\tilde{\theta}_{k}(X^{i}_{k})=\frac{\theta(X^{i}_{k})-\mu_{\theta_{k}}}{\sigma_{\theta_{k}}},$ (S4)
where $\mu_{\theta_{k}}$ and $\sigma_{\theta_{k}}$ denote the mean and standard deviation of $\theta$ across all walkers at time $t_{k}$. The rescaled score function $\tilde{\theta}_{k}$ is then used to compute the splitting function via
$V_{k}=C_{k}\tilde{\theta}_{k}.$ (S5)
Using this approach, we obtained results that were very similar to those achieved with the quantile-mapping step for large values of $N$ and better at small $N$.
S4.2 Scaling the constants $C_{k}$
With our choice of splitting function in Eq. (S5), some key parameters left to be tuned are the splitting constants $C_{k}$, which determine the rate at which walkers are duplicated at each resampling step.
A higher value of $C_{k}$ results in a larger number of walker clones, making the algorithm more selective. However, if $C_{k}$ is too large, it may lead to the extinction phenomenon described in Section S4.1, whereas if it is too small, the algorithm behaves similarly to direct numerical sampling by not sampling enough extreme events. It is often desirable not to use the same constant $C_{k}$ for all resampling steps. In our case, we are forecasting the future state of the trajectory to compute the score function, and the uncertainty in the forecast decreases as we approach the final time $t_{f}$. Therefore, we chose to set $C_{k}=Ce^{-\alpha(t_{f}-t_{k})}$, where $C>0$ is a constant and $\alpha>0$ is a decay factor. A similar approach is used in [14]. Both $C$ and $\alpha$ are hyperparameters of the algorithm. Using a non-exhaustive grid search, we found that setting $C=2.0$ and $\alpha=0.02$ yielded good results in terms of variance of the final estimators of rare event probabilities. Furthermore, we turned off resampling at the first three resampling times, $t_{1}$, $t_{2}$, and $t_{3}$, by setting $C_{1}=C_{2}=C_{3}=0$ to ensure that the walkers were sufficiently separated in phase space before actually applying a resampling.
S4.3 Choice of the resampling scheme
@@
S4.5 How we generate PlaSim ensembles
As described in the End Matter in the main text, for the duplication of walkers at each resampling step in the DMC algorithm to have an effect, it is necessary to perturb their initial conditions. Following [13], we introduce perturbations by adding a small noise to the coefficients of the spherical harmonics of the logarithm of the surface pressure field. Fig. S2 shows the time evolution of the standard deviation of 100-member PlaSim ensembles generated with different values of $\epsilon$. In our DMC implementation, we use a perturbation amplitude of $\epsilon=3\times 10^{-3}$. We tested several values of $\epsilon$ and found that this choice provides a good balance: it enhances the ability of the PlaSim ensembles to explore the phase space, while remaining small enough to avoid altering the final statistics of the target observable or introducing non-physical discontinuities in the trajectories. We anticipate that more sophisticated perturbation strategies, such as Bred vectors [66] or early splitting [17], could further improve the performance of the algorithm.
S5 Derivation of the computational speed-up factors
In the following, we derive two criteria to evaluate the computational gains provided by the AI+RES algorithm compared to DNS. We make the key assumption that the computational cost of running the AI emulator ensemble forecasts is negligible compared to that of running the physics-based model (in this case, the GCM). For reference, [27] report speed-ups of up to $\mathcal{O}(10^{4}\!-\!10^{5})$ and energy-consumption reductions of $\mathcal{O}(10^{4})$ for FourCastNet (AI emulator) relative to the IFS model (numerical weather prediction model). In our case, the cost comparison between the AI emulator and PlaSim is less favorable, as PlaSim is specifically designed to be computationally inexpensive. This is precisely what allows us to run a 50,000-member PlaSim ensemble within reasonable computational limits, enabling a robust validation of the methodology. Nevertheless, our analysis indicates that our emulator runs approximately 10 times faster on a single A100 GPU than PlaSim using 64 Intel Xeon Platinum 8160 CPU cores, even without any emulator variable or weight pruning or inference-specific optimizations. This does not undermine our assumption of negligible cost for the AI emulator, as the present study serves as a proof of concept.
S5.1 Sampling speed-up factor
$\text{SSUF}(a)=\frac{N_{DNS}(a)}{N_{RES}}=\frac{T_{a}n_{a}}{N_{RES}}$ (S7)
We also introduced a metric to quantify the diversity of the sampled trajectories (Section S7), demonstrating that the generated samples are not overly correlated.
S5.2 Variance speed-up factor
The other main objective of RES is to reduce the uncertainty in rare probability estimates. Following [14, 16], we also quantify speed-up based on the reduction in the variance of estimated rare event probabilities. Suppose we have used RES to obtain an estimate of the probability of exceedance $\hat{p}_{a}$ above the value $a$ with a variance $\sigma_{RES}(a)^{2}$ using a budget of $N_{RES}$ walkers. We can compare this with the variance of the DNS estimator (Eq. (S13)) as a function of $N$ and $a$, which is given by
$\sigma_{DNS}(a,N)^{2}=\frac{p_{a}(1-p_{a})}{N}$ (S8)
$N_{DNS}(a)\approx\frac{\hat{p}_{a}(1-\hat{p}_{a})}{\hat{\sigma}_{RES}(a)^{2}}.$ (S9)
We then define the Variance Speed-up Factor (VSUF) as
$\text{VSUF}(a)=\frac{N_{DNS}(a)}{N_{RES}}=\frac{p_{a}(1-p_{a})}{\hat{\sigma}_{RES}(a)^{2}N_{RES}}=\frac{T_{a}-1}{T_{a}^{2}\hat{\sigma}_{RES}(a)^{2}N_{RES}}$ (S10)
In practice, we estimate the variance of RES, $\hat{\sigma}_{RES}(a)^{2}$, empirically from only 10 independent realizations of the algorithm, as running a larger number of experiments is computationally expensive. Consequently, the estimate of $\hat{\sigma}_{RES}(a)^{2}$ itself is subject to sampling error. In Section S5.2.1, we also present alternative estimates of the variance using only a single realization of the algorithm.
S5.2.1 Computing error bars for the DMC algorithm
To compute the empirical variance estimates with the DMC algorithm, we performed 10 independent runs of the algorithm with the same hyperparameters and the same initial conditions. The same methodology is used in [12, 13, 15].
@@
The intuition behind these two forms is the following: the optimistic estimator considers each trajectory as an independent contribution, similar to importance sampling [68], while the pessimistic estimator aggregates all descendants of a given ancestor into a single effective data point. In practice, the effective number of independent samples lies between these two limiting cases. As there is no theoretical guarantee that one of these estimators is better than the other, we preferred to use the empirical variance computed from the 10 independent runs of the algorithm to compute variance speed-up factors (see Eq. (S10) and Fig. 3 in the main text). In Fig. S3, we verify that the empirical variance lies between the pessimistic and optimistic estimators, within statistical error.
@@
FIG. S3: Comparison of speed-up factors obtained via the empirical, the pessimistic (Eq. (S11)) and optimistic (Eq. (S12)) variance estimates. The experiments shown here are described in Table 1 in the main text, for Chicago (left panel) and France (right panel). The variance speed-up factor measures the reduction in the variance of return time estimates compared to direct numerical sampling (Eq. (S10)). The blue curve shows the empirical speed-up ratio obtained by running the AI+RES algorithm 10 times and computing the variance of the return time estimates across the 10 runs. The green curve shows the mean over the 10 optimistic variance estimates, each computed from a single realization of the algorithm. The orange curve shows the mean over the 10 pessimistic variance estimates, each computed from a single realization of the algorithm.
S6 Return period curves
$\mathbb{P}(A_{L}(t_{f})>a_{j})\approx\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}_{A_{L}(t_{f})>a_{j}}\left(Y_{t_{f}+L}^{i}\right)$ (S13)
For the RES algorithm, we use the formula (2) with the observable $\psi=\mathbbm{1}_{A_{L}(t_{f})>a_{j}}$ to compute the probability of exceedance. Namely,
$\mathbb{P}(A_{L}(t_{f})>a_{j})\approx\frac{\bar{w}_{K}}{N}\sum_{i=1}^{N}\mathbbm{1}_{A_{L}(t_{f})>a_{j}}\left(X_{t_{f}+L}^{i}\right)e^{-V_{K}\left(\hat{X}_{K}^{i}\right)}$ (S14)
S6.1 Interpretation of the return times
@@
FIG. S4: Most Recent Common Ancestor Distance (MRCAD) for AI+RES sampled trajectories. The figure shows results for France and Chicago with the setup given in Table 1 in the main text. The MRCAD is the average number of resampling steps since the current population last shared a common ancestor (max = 7, min = 1, with $K=6$ resampling steps). The solid lines show the mean MRCAD over 10 independent realizations of the AI+RES algorithm, with the shading indicating the $95\%$ confidence interval. A larger MRCAD indicates greater trajectory diversity. The AI+RES generated trajectories maintain substantial diversity, yielding a representative catalog of rare events.
@@
S8 Baseline methods
S8.1 Extreme value theory fits
$F_{(\sigma,\xi)}(z)=1-\left[1+\xi\frac{(z-u)}{\sigma}\right]^{-1/\xi}_{+},$ (S15)
To ensure a fair comparison in terms of sample size, the GPD distributions were fitted on datasets of the same size as the AI+RES experiments (namely $N=400$ in Fig. 2 of the main text), on 100 independent training datasets. The gray shading represents the 10th–90th percentile range of return-period curves obtained from the GPD fits across these datasets.
In the Peak-Over-Threshold method, the key hyperparameter to tune is the threshold $u$ above which the data is modeled. We determine $u$ using the heuristic of the Mean Residual Life plot ([69]), selecting a percentile that can be applied consistently across all datasets. Based on this criterion, we retain the 90th percentile of the data as the threshold. Parameters were estimated using a maximum likelihood approach with the scipy Python package ([70]).
S8.2 Using PlaSim itself to produce forecasts
@@
FIG. S5: Return-time curve comparison between AI+RES and PFS+RES. The setup matches that described in Table 1 in the main text, except that it uses $N=200$ walkers (instead of $N=400$), since larger values are computationally prohibitive for PFS+RES. Black dots are the empirical return periods obtained with a $50,000$-member control simulation. The red dots are from this control simulation but using the same computational budget as the AI+RES algorithm ($N=200$). The solid blue line shows the median return time curve produced by 10 independent realizations of the AI+RES algorithm with $N=200$ walkers and the blue-shaded area represents the 10th to 90th percentile range. The solid green line shows the median return time curve produced by 10 independent realizations of the PFS+RES algorithm with $N=200$ walkers and the green shaded area represents the 10th to 90th percentile range. The gray shaded area is obtained by fitting different GPD distributions with independent $N=200$ training datasets, and showing the 10th to 90th percentile range.
FIG. S6: Speed-up ratio for AI+RES and PFS+RES. The setup matches that described in Table 1, with France as the region of interest, except that it uses $N=200$ walkers (instead of $N=400$ in the main text), since larger values are computationally prohibitive for PFS+RES. Left: variance speed-up factor (Eq. (S10)). Right: sampling speed-up factor (Eq. (S7)), averaged over 10 independent algorithm realizations. Importantly, the speed-up factors are computed assuming that the cost of running the ensemble forecasts is negligible. This is certainly not the case for the PFS+RES algorithm. Instead, PFS+RES shows the speed-up factor that could be obtained if we had a perfect emulator.
@@
The Perfect-Forecast-System+RES (PFS+RES) approach is designed to provide an upper bound on algorithmic performance. As noted earlier, using the AI emulator introduces an imperfect approximation of the score function due to the emulator’s limited weather forecast skill, particularly beyond 10–15 days of lead time. Here, we propose performing the ensemble forecast directly with the PlaSim GCM, thereby obtaining a near-perfect forecast to use in the score function.
@@
We conducted experiments using the same setup as the AI+RES experiments, with parameters listed in Table 1 in the main text. The only difference is that the number of walkers was set to $N=200$ instead of $N=400$, as the latter is computationally prohibitive for PFS+RES.
@@
Fig. S5 shows the return-time curves for AI+RES and PFS+RES, demonstrating that PFS+RES can produce highly accurate return-period estimates with smaller variance than AI+RES. To quantify precisely how much we lose in terms of variance when moving from a perfect forecast system (PFS+RES) to an imperfect one (AI+RES), we compared the speed-up factors relative to DNS for both methods in Fig. S6. Importantly, these speed-up factors assume the cost of running ensemble forecasts is negligible. While this is not the case for PFS+RES, the comparison provides an indication of the potential performance achievable with near-perfect forecasts, under the assumption that forecast computations are inexpensive relative to the cost of running the physical model (here, PlaSim).
S9 Additional material
FIG. S7: Map of the regions of interest in this study. The left panel shows the region of France, and the right panel shows the region of Chicago. Each region is a box of $3\times 3$ pixels (for PlaSim resolution, each pixel is approximately 2.8 degrees). We chose to study mid-latitude heatwaves over France because previous studies using PlaSim or RES focused on this region [13, 45]. We also selected the Chicago region, as it is one of the hottest areas in summer in the PlaSim world, with the aim of sampling the most extreme events possible.
FIG. S8: Correlation between predicted and actual $A_{L}(t_{f})$ across forecast systems and regions. The blue curves correspond to forecasts made with the AI emulator within AI+RES for Chicago, while gold shows AI emulator results for France. Red curves represent forecasts made with the PlaSim GCM itself—referred to as the Perfect Forecast System (PFS) for the France region. Results are based on 10 independent realizations of the experiments in Table 1 in the main text. For PFS, the number of walkers is $N=200$ instead of $N=400$ for the AI+RES experiments, and the analysis is limited to France, for computational reasons. At each resampling step in AI+RES, score functions $\theta_{k}^{i}$ are computed from ensemble forecasts (with $M=100$ members) with the AI emulator for each PlaSim walker (Eq. (3) in the main text). Here, three resampling steps are used (at $t_{f}-t_{k}=10$, $5$, and 0 days before the event). We show the correlation between predicted (ensemble mean from the emulator) and actual (PlaSim realizations) $A_{L}(t_{f})$ for all events (left) and for events with return periods >100>100 years (right). For rare events, the emulator forecasts are notably less skillful at the earliest resampling time in France compared to Chicago, likely reflecting the stronger role of soil moisture in the former region (since soil moisture is absent from emulator inputs but evolved by PlaSim). This result explains the higher variance of rare-probability estimators in France compared to Chicago (left panel of Fig. 3 in the main text). Similarly, the smaller variances obtained for PFS-RES compared to AI-RES (Fig. S6), can also be attributed to an overall higher forecast skill of PFS both for typical and extreme events.
FIG. S9: T2m trajectories and histogram of the $A_{L}(t_{f})$ observable from a single AI+RES experiment over France. The setup matches the one described in Table 1. Left: solid blue lines show the evolution of daily T2m averaged over the region of interest for all $N=400$ walkers. At each resampling time (vertical dashed lines), ensemble forecasts with the emulator are run for each PlaSim walker until the end of the simulation, and the most promising are duplicated based on these forecasts (Eq. (3)). In particular, at the very bottom of the figure at the resampling time of 07/27, we observe that a large number of clones are drawn for a walker that exhibited a low T2m value at the time of resampling, but whose temperature rapidly increased thereafter, illustrating the emulator’s ability to anticipate this event. The red dashed line at $t_{f}$ marks the start of the event of interest on 08/01. Right: histograms of the distribution of $A_{L}(t_{f})$ (Eq. (1) with $L=7$ days) for direct numerical simulation (DNS, gray) and AI+RES (blue). The shift between them illustrates the algorithm’s importance sampling effect.
FIG. S10: Histogram of the $A_{L}(t_{f})$ observable in the training dataset of the AI emulator for the France (gold) and Chicago (blue) regions with $L=7$ days and $t_{f}$ = August 1. The maximum value seen during training by the AI emulator is $33.04\text{\textdegree C}$ in France and $44.28\text{\textdegree C}$ in Chicago. As seen in the AI-DNS results (cyan dots) in Fig. 2, the AI emulator is able to simulate events much more extreme than what was seen during the training.
FIG. S11: Scaling of the relative error on the estimation of the probability $p=0.001$ as a function of the number of walkers NN. The setup matches the one described in Table 1 (except the varying NN), with France as the region of interest. The relative error for the probability $p$ is defined as $\mathbb{R}\mathbb{E}:=\frac{\sqrt{\mathbb{V}\left(\hat{p}\right)}}{\mathbb{E}\left(\hat{p}\right)}$. The slope of the DNS curve (red line) is 1/2-1/2, as can be derived from Eq. (S8). For AI+RES, we observe that increasing NN yields a greater reduction in relative standard deviation when the number of walkers is low (between 50 and 200), while for larger NN, the improvement becomes comparable to that of DNS.
FIG. S12: Return-time curves of $A_{L}(t_{f})$ for $L=1,3,5$ and 77 days (Eq. (1)) and France as the region of interest. The results shown here are from the same experiments presented in Table 1. In particular, while the score function used here is the 7-day average surface temperature over the region of interest, the trajectories sampled by the AI+RES algorithm allow us to accurately estimate the return period of shorter-duration events. The darker dots are the empirical return periods obtained with a $50,000$-member control simulation. The solid lighter lines show the mean of return time curves produced by 10 independent realizations of the AI+RES algorithm with $N=400$ walkers. The light shaded areas represent $95\%$ confidence intervals.